OpenStudy (anonymous):

ken is making a large batch of his cookies and will use 3 cups of brown sugar based on this information what is the total number of cupes of flour will he need to use

OpenStudy (anonymous):

I don't care what kind of cookies Ken is making. I don't even care that this isn't a biology question. If you want help, then you need to show us the original recipe. I can show you how to cross multiply. Because that's what you need to do, to solve this problem. For example, if the original recipe calls for 2 cups of flour and 1 cup of sugar, then this recipe will need 6 cups of flour, to go with those 3 cups of sugar. Show us the original recipe.

OpenStudy (anonymous):

• 1 cup butter • ¾ cups sugar • ¾ cups brown sugar • 2 cups flour • 2 large eggs • 1 teaspoon vanilla extract • 2 cups semi-sweet chocolate chips • 1 cup chopped pecans

OpenStudy (anonymous):

@livenaldo @coolkidtati I'm gonna assume that this is the original recipe, and that you two are in the same class. ~~~~~~~~~~~~~ Okay. The original recipe tells us that for 2 cups of FLOUR, we need 3/4 cup of (brown) SUGAR. We need to put that into a fraction, with FLOUR on TOP, and SUGAR on the BOTTOM. But first, let's change 3/4 into a decimal number, to make it easier for us. 3/4 is equal to 0.75. Yes? Okay. So now let's make that fraction. Again, 2 cups of FLOUR for 0.75 cup of brown SUGAR. \[\frac{ 2 }{ 0.75 }\] Now, that fraction is only for the original recipe. FLOUR (2) on TOP, and SUGAR (0.75) on the BOTTOM, you see? Hold onto that fraction, because that's only the fraction for the original recipe. Now we need to make a fraction for the recipe in the question. The question asks us... "How many cups of FLOUR would be need, if we had 3 cups of brown SUGAR?" Let's set up another fraction. FLOUR on TOP, SUGAR on the BOTTOM. We don't know how much FLOUR we'll need, so, to represent flour, we'll make an X. But we know that we'll need 3 cups of SUGAR, yes? So here's the fraction for the recipe in the question: \[\frac{ x }{ 3 }\] ~~~~~~~~~~~~~ Now, we have both the fraction for the original recipe, and the fraction for the recipe in the question. Let's put them side by side, with an equal sign. Like this: \[\frac{ 2 }{ 0.75 } = \frac{ x }{ 3 }\] Now, we need to solve for how much flour we'll need, for 3 cups of brown SUGAR. We need to solve for X. And we do that by CROSS MULTIPLYING. There are two steps: 1. MULTIPLY the 2, and the 3. \[2\times3 = 6\] 2. Take 6, and DIVIDE it by 0.75 \[6\div0.75 = 8\] That means X = 8. So, let's look at the completed fractions again, side by side: \[\frac{ 2 }{ 0.75 } = \frac{ 8 }{ 3 }\] Remember, FLOUR was on TOP, and brown SUGAR was on the BOTTOM. Look at those two fractions. If we have 2 cups of FLOUR, for 0.75 cup of brown SUGAR, then we would need 8 cups of FLOUR for 3 cups of brown SUGAR. Do you get it? And all we had to do was cross multiple, in two easy steps.